OpenStudy (anonymous):

Write an expression for the ratio of the perimeter to the area of a square with sides of 8x

OpenStudy (anonymous):

Perimeter = 4 times side lenght = 4 * 8x = 32x Area = side * side = 8x * 8x = 64x^2 Ratio = Perimeter/Area = 32x / 64x^2 = 1/(2x)

OpenStudy (unklerhaukus):

\[\frac P A =\frac{4s}{s^2}=\frac{4}{s}\]

OpenStudy (unklerhaukus):

now let \[s=8x\]

OpenStudy (anonymous):

Would that be the ratio of perimeter to area?

OpenStudy (unklerhaukus):

\[P/A=\frac{4}{s}=\frac{4}{8x}=\frac{1}{2x}\]

OpenStudy (unklerhaukus):

i am using P for perimeter and A of area the ratio of P to A is P over A

OpenStudy (anonymous):

Thanks